OpenStudy (theedwardsfamily):

Which circles have a circumference of 24\(pi)\ units? Choose all answers that are correct. A. a circle with a radius of 6 units B. a circle with a radius of 12 units C. a circle with a diameter of 12 units D. a circle with a diameter of 24 units

OpenStudy (epoweritheta):

calculate the circum for each of the option s . tell me for optn A ?>?

OpenStudy (theedwardsfamily):

6*3.14 ?

OpenStudy (epoweritheta):

remember , c= 2 * pi * r

OpenStudy (theedwardsfamily):

idk how to do

OpenStudy (epoweritheta):

u dont know to multiply ????? 2* 3.14*6 ?

OpenStudy (theedwardsfamily):

like order of operations?

OpenStudy (theedwardsfamily):

i got 37.68

OpenStudy (theedwardsfamily):

@AlexandervonHumboldt2

OpenStudy (epoweritheta):

ya correct

OpenStudy (epoweritheta):

See 24 pi units = 2* 12 * pi

OpenStudy (theedwardsfamily):

75.36

OpenStudy (theedwardsfamily):

@epoweritheta

OpenStudy (epoweritheta):

okay check other options then

OpenStudy (theedwardsfamily):

they add up

OpenStudy (theedwardsfamily):

@jim_thompson5910

OpenStudy (theedwardsfamily):

@kidrah69

jimthompson5910 (jim_thompson5910):

Circumference C = pi*d or C = 2*pi*r

jimthompson5910 (jim_thompson5910):

since the circumference is 24pi, we can use C = pi*d to determine that the diameter is 24 if the diameter is 24, then the radius must be 24/2 = 12